CH3COOH(aq)+NaOH(aq) ---> CH3COONa(aq)+H2O(l)
10ml unknown concetration ethanoic acid CH3COOH was titrated by 0,1mol/l concetration NaOH solution. While looking at this data:
(If NaOH volume is 0ml=4pH;10ml=4,5pH;15ml=5,2pH;20ml=5,6pH;25ml=6pH;28ml=6,5pH;30ml=6-7pH;32ml=12,3pH;34ml=12,6pH;40ml=13pH)
Please Find ethanoic acid's concetration. Write a consistent answer.
1
Expert's answer
2019-11-26T06:54:10-0500

Solution.

Based on the data on the pH change depending on the added alkali, we see that the equivalence point is at a volume of 30 ml.

Write down the law of equivalents:

"C1 \times V1 = C2 \times V2"

"C(CH3COOH) = \frac{C(NaOH) \times V(NaOH)}{V(CH3COOH)}"

C(CH3COOH) = 0.3 mol/L

Answer:

C(CH3COOH) = 0.3 mol/L